OpenStudy (anonymous):

Akua used a globe and a measuring tape to find the distance between two cities. The two cities on the globe were 3 inches apart. The scale of the globe was 1 inch : 656 miles. What is the actual distance between the two cities?

OpenStudy (anonymous):

A 1,312 miles B 1,968 miles C 659 miles

OpenStudy (anonymous):

Omgg yay jimmmm ur on

jimthompson5910 (jim_thompson5910):

1/656 = 3/x 1*x = 656*3 ... cross multiply x = ???